Find: (i) 0.2 × 6 (ii) 8 × 4.6 nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Find:
(i) 0.2 × 6

(ii) 8 × 4.6

 

nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(i) 0.2 × 6

Solution:- We have,

= (2/10) × 6

= (12/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place. Then,

= 1.2

(ii) 8 × 4.6

Solution:- We have,

= (8) × (46/10)

= (368/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place. Then,

= 36.8


Thanks